1. A food safety guideline is that the mercury in fish should be below 1 part per million (ppm). Listed below are the amounts of mercury (ppm) found in tuna sushi sampled at different stores in a major city. Find the margin error for 95% confidence interval and construct a 95% confidence interval estimate of the mean amount of mercury in the population. Does it appear that there is too much mercury in tuna sushi? Mecury(ppm) .50 .079 .0.09 .89 1.29 .59 .91.

2. An IQ test is designed so that the mean is 100 and the standard deviation is 23 for the population of normal adults. Find the sample size necessary to estimate the mean IQ score of statistics students such that it can be said with 90% confidence that the sample mean is within 6 IQ points of the true mean. Assume that =23 and determine the required sample size using technology.

pls do it yourself, thanks!
Answer #1

(1)

n= 7
sample mean= 0.621
s= 0.446

(1-alpha)*100% confidence interval for population mean=sample mean±t(alpha/2,n-1)*s/sqrt(n)

95% confidence interval for population mean=mean±t(0.05/2, n-1)*s/sqrt(n)=(0.209,1.034)

t-value margin of error lower limit upper limit
95% confidence interval 2.447 0.413 0.209 1.034

(2)

let sample size=n then  

with (1-alpha)*100% confidence margin of error=z(alpha/2)*sd/sqrt(n)

with 90% confidence margion of error=z(0.1/2)*sd/sqrt(n)

or, 6=1.645*23/sqrt(n)

or, sqrt(n)=1.645*23/6=6.31

n=39.82 ( next whole number is 40)

answer =40